Draw the graph of the function f :R---R :f(x)= x^2-6x+5. Obtain the coordinates of the point at which the graph meets the x and y axes.
1
Expert's answer
2019-03-20T09:58:42-0400

Solution:

If the function

"f:R\mapsto{R}"

and

"f(x)=x^2-6x+5"

then the graph of the function looks like this



where


"(1;0), (5;0)"

are the points at which the graph meets the x axis because

"x^2-6x+5=0,D=36-20=16,"

"x_1=\frac{6-\sqrt{16}}{2}=1,
x_2=\frac{6+\sqrt{16}}{2}=5;"

and


"(0;5)"

is the points at which the graph meets the y axis because


"0^2-{6}\cdot{0}+5=5."

Answer:


"(1;0),(5;0),(0;5)"

are the points at which graph meets the x and y axes.